OpenStudy (anonymous):

Find the volume of a sphere with radius 6 ft leave answer in terms of pi.

OpenStudy (anonymous):

4/3πr3

OpenStudy (anonymous):

I believe it is \[288\pi \] The equation of a sphere is \[\frac{ 4 }{ 3 }\times \pi \times R ^{3}\] Just plug in 6ft into R for the radius.